OpenStudy (anonymous):

compute the first derivates of the following function y=ln((1+3x)^4/(sq rt(2x+2))

zepdrix (zepdrix):

Everything is inside the log? Hmmm I guess we can take advantage of some of our rules of logarithms :) to break it down, because we don't want to apply the chain rule, quotient rule, chain chain. That just sounds like a pain in the butt. \[\log(ab)=loga+logb\]\[\log(\frac{ b }{ c })=logb-logc\] Can we use these rules somehow? :)